OpenStudy (anonymous):

Let \[ 3 x + 4 y = 0\\ 1 + 5 x + 12 y = 0\\ \] be two intersecting lines. Le I be the point of intersection of these two lines. Find the equations of the two bisectors lines of the two angles formed by the the two above points at the point I.

OpenStudy (anonymous):

@rational

terenzreignz (terenzreignz):

Interesting.... perhaps putting them in y=mx+b form might lead me somewhere :D \[\Large y = -\frac{4}{3}x\]\[\Large y = -\frac{5}{12}x-1\]

OpenStudy (anonymous):

@terenzreignz , I got different equations.

terenzreignz (terenzreignz):

oops... \[\Large y = -\frac{5}{12}x-\frac1{12}\]

OpenStudy (anonymous):

OK. This is the equation of one of the line and not the answer.

terenzreignz (terenzreignz):

mhmm :3 and now... I'm thinking.

OpenStudy (anonymous):

Here is a graph of the lines and their two bisectors. The red and green are the the two lines, the others are their bisectors.

terenzreignz (terenzreignz):

Do you already have the answer? I have a feeling they'll be looking far from pretty....

OpenStudy (anonymous):

Not really, it involves only rational numbers.

terenzreignz (terenzreignz):

That's interesting... and encouraging... let me see :3

terenzreignz (terenzreignz):

I guess just take the mean values of the slopes?

OpenStudy (anonymous):

Look at the following for a hint: http://openstudy.com/users/eliassaab#/updates/52d8eefee4b05ee05437b571

terenzreignz (terenzreignz):

scratch the last one...

terenzreignz (terenzreignz):

Crud... I hate it when I go wrong... for messing up details... it should have been \[\Large y = -\frac34x\]and\[\Large y=-\frac{5}{12}x-1\]

terenzreignz (terenzreignz):

Then, translate: \[\Large \bar x := x- \frac14\]\[\Large \bar y :=y+\frac3{16}\]

terenzreignz (terenzreignz):

Equations become \[\Large \bar y = -\frac34 \bar x\]\[\Large \bar y = -\frac5{12}\bar x\]

terenzreignz (terenzreignz):

Find a point on the first line which is 1 unit away from the solution (centre). \[\Large \sqrt{\bar x^2 + \frac{9}{16}\bar x^2}=\frac54\bar x = 1\implies \bar x= \frac45; \bar y = -\frac35\] Similarly, on the second line, a point on the line which is 1 unit away from the solution is \[\Large \bar x = \frac{12}{13};\bar y = -\frac{5}{13}\]

terenzreignz (terenzreignz):

The midpoint of these two points is \[\Large \bar x =\frac{56}{65};\bar y =-\frac{32}{65}\]

terenzreignz (terenzreignz):

This and the point \[\Large \bar x = 0;\bar y = 0\] determine a line which is an angle bisector of the original two lines: That line is given by\[\Large \bar y=-\frac{4}{7}\bar x\] The other angle bisector is the line through the centre which is perpendicular to this line: \[\Large \bar y = \frac74x\]

terenzreignz (terenzreignz):

Reversing the translation and then simplifying, the lines are: \[\Large y =-\frac47x-\frac5{112}\]\[\Large y=\frac74x-\frac58\] @eliassaab ?

terenzreignz (terenzreignz):

Oh, and I messed up my earlier correction (-_- what is wrong with me today) \[\Large y = -\frac34x\]\[\Large y = -\frac5{12}\color{red}{(x+1)}\] anyway, that aside, what do you think @eliassaab ?

OpenStudy (anonymous):

@terenzreignz your equations are correct. That is what I got

OpenStudy (rational):

using the hint from previous problem : every point on angle bisector is equi distant from both the arms say, P(x, y) be any point on the bisector, then this satisfies :- \(\large \frac{3x+4y}{\sqrt{3^2+4^2}} = \pm \frac{1+5x+12y}{\sqrt{5^2+12^2}}\) \(\large \frac{3x+4y}{5} = \pm \frac{1+5x+12y}{13}\) \(\large y = \frac{1}{8} (14x-5)\) \(\large y = \frac{1}{112}(-64x-5)\)

OpenStudy (anonymous):

@rational, this is how I did it.